OpenStudy (anonymous):

What value for y makes this equation true? 3 • y = (3 • 40) + (3 • 8)

OpenStudy (anonymous):

@A_Graff

OpenStudy (anonymous):

helooooo

OpenStudy (anonymous):

come on

OpenStudy (anonymous):

I believe its 40

OpenStudy (anonymous):

Let's work this out. (3*40)=120, correct? (3*8)=24, right? Add them together to get 144. What multiplied by 3 gives us 144?

OpenStudy (anonymous):

3*48 = 144 @mossyfish

OpenStudy (anonymous):

That's right! Good job :)

OpenStudy (anonymous):

Will u help with another quastion?

OpenStudy (anonymous):

@mossyfish

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Which is another way to write the expression 25 • (r + 6)?

OpenStudy (anonymous):

What are your options? There are many ways to rewrite that expression.

OpenStudy (anonymous):

A. (25 • r) + (25 • 6) B. (25 • r) + 6 C. 25 • r + 6 D. 25 • r • 6

OpenStudy (anonymous):

Let's try expanding this. 25(r+6) is equal to 25*r and 25*6, correct? How would you write this?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Let's say that r=1. 25*1*6 is equal to 150. 25*(1+6) is 175. So, substituting 1 for r, which of those is equal to 175?

OpenStudy (anonymous):

so C

OpenStudy (anonymous):

No... 25*1+6 is 31.

OpenStudy (anonymous):

Here, I think that you need to review the distributive property https://www.khanacademy.org/math/cc-sixth-grade-math/cc-6th-factors-and-multiples/cc-6th-distributive-property/v/the-distributive-property

OpenStudy (anonymous):

i am so confused will u just give me a answer

OpenStudy (anonymous):

Can you show me your work? Maybe if I know where you're going wrong I can help you

OpenStudy (anonymous):

I dont have much time i am timed and i really need help\

OpenStudy (anonymous):

I'm not just going to give you the answer. We ruled out D and C. Now, substitute 1 for r and solve A and B. Which one of those gives you 175?

OpenStudy (anonymous):

Its B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It's not B. Let's expand the original problem. 25(r+6). We have to FOIL. 25*r and 25*6. This is not equal to (25*r)+6 (B). We are multiplying 6, not adding it. This gives us A, (25*r)+(25*6). I think that you should ask your teacher to explain this to you